OpenStudy (anonymous):

Classify √5 A. Whole Number B. Integer C. Rational Number D. Irrational Number

OpenStudy (bahrom7893):

irrational number if im not mistaken

OpenStudy (bahrom7893):

am i right?

OpenStudy (jamesj):

It's certainly not a whole number of integer, because 2^2 = 4 and 3^2 = 9. And it also turns out it's not a rational number, although that's harder to prove. But yes, you're right: \( \sqrt{5} \) is irrational.

OpenStudy (anonymous):

my guess is you are just supposed to know that \[\sqrt{a}\] is irrational unless \[a\] is a perfect square (like 16 or 25)

OpenStudy (anonymous):

by "know" i mean 'believe'